A wind turbine was constructed and placed in service by an energy company. Its cost basis is $83,000, and it has an estimated MV of $3,200 at the end of an estimated useful life of 10 years. Compute the book value (BV) at the end of the array's eighth year of life when using the DB method with switchover to SL and a DB ratio of 200%. Express your answer in terms of dollars and cents, rounded to the nearest cent (e.g., 1234.56).
Answer #1

Annual Depreciation rate = (100%/useful life) = (100%/10) = 10%

DB ratio of 200% is given. Under DB method of depreciation, depreciation is charged at the double rate than the annual depreciation rate. so,

Annual depreciation under DB method = 10%*2 = 20%

estimated MV of wind turbine is $3,200 at the end of an estimated useful life of 10 years.

depreciable value = cost basis - estimated MV or salvage value = $83,000 - $3,200 = $79,800

Years Beginning depreciable value Depreciation Ending book value
1 $79,800.00 $15,960.00 $63,840.00
2 $63,840.00 $12,768.00 $51,072.00
3 $51,072.00 $10,214.40 $40,857.60
4 $40,857.60 $8,171.52 $32,686.08
5 $32,686.08 $6,537.22 $26,148.86
6 $26,148.86 $5,229.77 $20,919.09
7 $20,919.09 $4,183.82 $16,735.27
8 $16,735.27 $3,347.05 $13,388.22

the book value (BV) at the end of the array's eighth year of life when using the DB method is $13,388.22.

Calculations

2 Years Beginning depreciable value 79800 ED3 D4 Depreciation =B3*20% =B4*20% =B5*20% =B6*20% =B7*20% =B8*20% =B9*20% =B10*20